OpenStudy (lemanual):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits.

OpenStudy (anonymous):

Just multiply 6.25 6 times. Remember a square has 6 congruent sides.

OpenStudy (lemanual):

Ty :)

OpenStudy (anonymous):

No problem! :)

OpenStudy (lemanual):

If I'm right is it 59604.64477539063

OpenStudy (solomonzelman):

\(\large\color{slate}{ 25^2=625 }\) and same way is, \(\large\color{slate}{ 2.5^2=6.25 }\)

OpenStudy (solomonzelman):

the area of square is \(\large\color{slate}{ A=(side)^2 }\)

OpenStudy (lemanual):

So it's 2.5 2 I don't know how to put the 2 at the top

OpenStudy (solomonzelman):

oh, the side is 6.25 I thought it is the entire area

OpenStudy (solomonzelman):

\(\large\color{slate}{ 625^2=390625 }\) and so, \(\large\color{slate}{ 6.25^2=39.0625 }\)

OpenStudy (solomonzelman):



OpenStudy (lemanual):

Ty So Much :)

OpenStudy (solomonzelman):